OpenStudy (anonymous):

How would you write y-7=-3/4(x+5) from point slope form to slope intercept form ?? -3/4 is a fraction in the equation btw. Thank you

OpenStudy (texaschic101):

y - 7 = -3/4(x + 5) -- distribute through the parenthesis y - 7 = -3/4x - 15/4 -- add 7 to both sides y = -3/4x - 15/4 + 7 -- common denominator is 4 y = -3/4x - 15/4 + 28/4 y = -3/4x + 13/4 <== slope intercept form any questions ? Basically, all you do is solve for y by moving everything else over to the other side.